The table shows several value of the function f(x)=-x^3+x^2-x+2. Complete the missing values in the table, and then use these values and the imtermediate value theorem to determine (an) interval(s) where the function must have zero.
Table:
X |-2|-1| 0| 1| 2|
y |16| | | | |
I initially had 8, 0, -8, -4 in the blank spots of teh table for y but I am not sure if that is correct.

Answer by stanbon(48535)

f(x)=-x^3+x^2-x+2.
X= |-2|-1| 0| 1| 2|
f(-2)= 16
f(-1)=-(-1)^3 +(-1)^2-(-1)+2
=8+1+1+2=12
f(0)=2
f(1)=-1+1-1+2=1
f(2)=-8+4-2+2=-4
Cheers,
Stan H.